# Asymptotics of high-codimensional area-minimizing currents in hyperbolic space

### v1

@@ 1. Introduction
Let $T$ be an area-minimizing locally rectifiable $n$-current in $\mathbb{H}^{\mathfrak{m}}$ that is asymptotic to $\Gamma$, which is a closed (n1)(n-1)-dimensional $C^{1,\alpha}$ submanifold of $\mathbb{R}^{\mathfrak{m}-1}\times\{0\}$ for some $0<\alpha\leqslant 1$. Let $H^{+},G_{R},B_{R}^{+}$ be defined as in (1.3), (1.6), and (1.7), respectively, with the coordinate condition (1.5) satisfied on $H^{+}$. There exists a constant $R>0$ such that the following hold:
@@
We note that Assumption 1.3 is weaker than Assumption 1.1. By the Constancy Theorem (4.1.7, [4]), there exists a small constant $\rho_{\Gamma}>0$ depending on $\Gamma$ such that the validity of (1.8) for a single $r\in(0,\rho_{\Gamma})$ is equivalent to its validity for all $r\in(0,\rho_{\Gamma})$.
@@
in (1.9) grows doubly exponentially with respect to the geodesic distance function on $\mathbb{H}^{\mathfrak{m}}$.
@@
We next introduce the corresponding system of equations and proceed to investigate its regularity theory; in general, however, the analysis of such regularity theory poses substantial challenges owing to the absence of a maximum principle. Examples demonstrate that solutions of the system of minimal surface equations may fail to exist, be non-unique, or lack stability; for instance, even when the domain is a 4-dimensional ball and the boundary values are analytic, Lipschitz solutions to the minimal surface system generally do not exist. While Lawson and Osserman [22] established the existence of solutions to the minimal surface system when the domain is 2-dimensional, they also constructed examples showing that such solutions are generally neither unique nor stable. Within the hyperbolic setting, we have rigorously derived a particularly elegant system of equations governing minimal surfaces near $\Gamma$, which we present as follows.
@@
Let $T$ be an area-minimizing locally rectifiable current in $\mathbb{H}^{\mathfrak{m}}$. Suppose that near $Q\in\Gamma$, $T$ can be represented as the graph of the mapping
@@
By analyzing the regularity of $\mathbf{u}$, we establish the following theorems.
@@ 3. From Local Rectifiability to $C^{1,\alpha}$ Regularity
In $B_{\mathbb{H}^{\mathfrak{m}}}(P,2)$, this integrand is (with respect to the Euclidean metric) elliptic with ellipticity bound $e^{-2n}$.
@@
Let $T$ be an area-minimizing locally rectifiable $n$-current in $\mathbb{H}^{\mathfrak{m}}$, and let $P$ be a point in $H^{+}$ as defined in (1.3). Assume that $T$ is a normal current in the standard Euclidean metric with $\partial T=[\Gamma]$, where $\Gamma$ is a $C^{1,\alpha}$ submanifold of the hyperplane at infinity. Then there exists a constant $\rho_{\Gamma}>0$ such that if
$x^{\mathfrak{m}}(P)<\tfrac{1}{2}\rho_{\Gamma},$ (3.18)
we have
@@
where the constant $C=C(\mathfrak{m},n,\Gamma)$ is independent of the choice of $P$.
@@
$\leqslant e^{-\frac{1}{c_{0}\delta}}M_{\mathbb{H}^{\mathfrak{m}}}\bigl((\Phi_{\#}T)\mathbin{\llcorner}B_{\mathbb{H}^{\mathfrak{m}}}(\Phi(P),2)\bigr)+C(\mathfrak{m},n)$
$=e^{-\frac{1}{c_{0}\delta}}M_{\mathbb{H}^{\mathfrak{m}}}\bigl(T\mathbin{\llcorner}B_{\mathbb{H}^{\mathfrak{m}}}(P,2)\bigr)+C(\mathfrak{m},n)$
$\leqslant e^{-\frac{1}{c_{0}\delta}}e^{2n}\cdot\bigl(x^{\mathfrak{m}}(P)\bigr)^{-n}M_{\mathbb{R}^{\mathfrak{m}}}(T)+C(\mathfrak{m},n)$
@@
where $C$ in (3.31) depends on $\mathfrak{m},n$ and $\Gamma$. ∎
By the proof of Theorem 3.3, we in fact establish the following lemma.
@@
Let $T$ be an area-minimizing locally rectifiable $n$-current in $\mathbb{H}^{\mathfrak{m}}$ that is asymptotic to $\Gamma\in C^{1,\alpha}$, satisfying
$\partial T\mathbin{\llcorner}B_{\mathbb{H}^{\mathfrak{m}}}(P,2)=0$ (3.32)
and (3.15). Let $\rho_{\Gamma}$ be defined as in Section 2. If in addition, for some given constant $c_{0}>0$ and for some $P\in\mathbb{H}^{\mathfrak{m}}$ with $x^{\mathfrak{m}}(P)<\tfrac{1}{2}\rho_{\Gamma}$, the bound
@@
By Theorem 3.3, for every $P\in\mathbb{H}^{\mathfrak{m}}$ with $x^{\mathfrak{m}}(P)<\tfrac{1}{2}\rho_{\Gamma}$, we have
$M_{\mathbb{H}^{\mathfrak{m}}}\bigl(T\mathbin{\llcorner}B_{\mathbb{H}^{\mathfrak{m}}}(P,1)\bigr)<C(\mathfrak{m},n,\Gamma)\bigl(M_{\mathbb{R}^{\mathfrak{m}}}(T)+1\bigr).$ (3.35)
@@
$\|\textbf{u}\|_{C^{1,1}}\leqslant\varepsilon\bigl(x^{\mathfrak{m}}(P),\mathfrak{m},n,\Gamma\bigr),$ (3.37)
where
@@
Theorem 1.4 follows by the same argument.
@@ 4. System of Minimal Surface Equations
(1.6). Setting $y^{1}=x^{1},\cdots,y^{n-1}=x^{n-1},y^{n}=x^{\mathfrak{m}}$ as in (1.11), we may rewrite this graph mapping as
@@
If all mean curvatures $H_{s}$ vanish, then $\Delta_{g_{M}}\mathbf{F}=0$. To establish this, we start with the covariant derivative expansion of $\mathbf{F}_{,ij}$:
@@
$\nabla_{\frac{\partial}{\partial y^{l}}}\frac{\partial}{\partial y^{k}}=-\frac{1}{y^{n}}\left(\delta_{l,n}\frac{\partial}{\partial y^{k}}+\delta_{k,n}\frac{\partial}{\partial y^{l}}-\delta_{l,k}\frac{\partial}{\partial y^{n}}\right).$ (4.12)
@@
$g_{M}^{ij}\frac{\partial\mathbf{F}_{l}}{\partial y^{i}}\frac{\partial\mathbf{F}_{k}}{\partial y^{j}}\cdot\nabla_{\frac{\partial}{\partial y^{l}}}\frac{\partial}{\partial y^{k}}=\begin{cases}-\frac{1}{y^{n}}g_{M}^{nn}\frac{\partial}{\partial y^{n}},&l,k=n;\\ -\frac{1}{y^{n}}\frac{\partial u_{l}}{\partial y^{i}}g_{M}^{in}\frac{\partial}{\partial y^{l}},&l=1,\cdots,n-1\text{ and }k=n;\\ \frac{1}{y^{n}}\frac{\partial u_{l}}{\partial y^{i}}\frac{\partial u_{k}}{\partial y^{j}}\delta_{l,k}g_{M}^{ij}\frac{\partial}{\partial y^{n}},&l,k=1,\cdots,n-1.\end{cases}$ (4.13)
@@
If the graph of $\mathbf{u}$ in the form given by (4.3), defined over a domain $\Omega\subseteq H^{+}$, is an $n$-dimensional minimal submanifold of $\mathbb{H}^{\mathfrak{m}}$ (i.e., all mean curvatures vanish), then the PDE system (4.15) holds.
We may also derive the system (4.15) via the method of variations.
Consider an arbitrary smooth map defined on $G_{r}$, given in the form
@@ 5. Higher Order Regularity Theorems
This equation is uniformly elliptic after rescaling in a neighborhood of any point $P\in G_{r/2}$. Precisely, under the rescaled coordinate transformation

### v2

@@ 1. Introduction
Let $T$ be an area-minimizing locally rectifiable $n$-current in $\mathbb{H}^{\mathfrak{m}}$ that is asymptotic to $\Gamma$, which is a closed (n1)(n-1)-dimensional $C^{1}$ submanifold of $\mathbb{R}^{\mathfrak{m}-1}\times\{0\}$ and
$\Gamma\cap\overline{G_{R}}\in C^{1,\alpha},$ (1.8)
for some $0<\alpha\leqslant 1$. Let $H^{+},G_{R},B_{R}^{+}$ be defined as in (1.3), (1.6), and (1.7), respectively, with the coordinate condition (1.5) satisfied on $H^{+}$. There exists a constant $R>0$ such that the following hold:
@@
We note that Assumption 1.3 is weaker than Assumption 1.1. By the Constancy Theorem (4.1.7, [4]), there exists a small constant $\rho_{\Gamma}>0$ depending on $\Gamma$ such that the validity of (1.9) for a single $r\in(0,\rho_{\Gamma})$ is equivalent to its validity for all $r\in(0,\rho_{\Gamma})$. Equation (1.9) enforces the multiplicity one condition; this serves to exclude singularities, such as branch points, which occur in higher multiplicity scenarios.
@@
in (1.10) grows doubly exponentially with respect to the geodesic distance function on $\mathbb{H}^{\mathfrak{m}}$. A key insight in our work is that the doubly exponential mass growth condition implies a uniform upper bound for the local mass near the asymptotic boundary.
@@
We next introduce the corresponding system of equations and proceed to investigate its regularity theory; in general, however, the analysis of such regularity theory poses substantial challenges owing to the absence of a maximum principle. Examples demonstrate that solutions of the system of minimal surface equations may fail to exist, be non-unique, or lack stability; for instance, even when the domain is a 4-dimensional ball and the boundary values are analytic, Lipschitz solutions to the minimal surface system generally do not exist. While Lawson and Osserman [22] established the existence of solutions to the minimal surface system when the domain is 2-dimensional, they also constructed examples showing that such solutions are generally neither unique nor stable. Within the hyperbolic setting, we have rigorously derived an elegant system of equations governing minimal surfaces near $\Gamma$, which we present as follows.
@@
Let $T$ be an area-minimizing locally rectifiable current in $\mathbb{H}^{\mathfrak{m}}$. Suppose that near $Q\in\Gamma$, $T$ can be represented as the graph of a $C^{1}$ mapping
@@
The regularity analysis of $\mathbf{u}$ yields the following theorems, which settle the PDE aspects of the asymptotics for area-minimizing locally rectifiable currents in hyperbolic space.
@@ 3. From Local Rectifiability to $C^{1,\alpha}$ Regularity
In $B_{\mathbb{H}^{\mathfrak{m}}}(P,2)$, this integrand is (with respect to the Euclidean metric) elliptic with an ellipticity bound $e^{-2n}$.
@@
Let $\Gamma$ be a $C^{1,\alpha}$ submanifold of the hyperplane at infinity, let $\rho_{\Gamma}$ be defined as in Section 2 and let $P$ be a point in $G_{r}$ with
$r<\tfrac{1}{2}\rho_{\Gamma}.$ (3.18)
Assume that $T$ is an area-minimizing locally rectifiable $n$-current in $\mathbb{H}^{\mathfrak{m}}$, and $T$ is a normal current in the standard Euclidean metric with $\partial T=[\Gamma]$. Then
@@
where the constant $C=C(\mathfrak{m},\Gamma)$ is independent of the choice of $P$.
@@
$\leqslant e^{-\frac{1}{c_{0}\delta}}M_{\mathbb{H}^{\mathfrak{m}}}\bigl((\Phi_{\#}T)\mathbin{\llcorner}B_{\mathbb{H}^{\mathfrak{m}}}(\Phi(P),2)\bigr)+C(\mathfrak{m})$
$=e^{-\frac{1}{c_{0}\delta}}M_{\mathbb{H}^{\mathfrak{m}}}\bigl(T\mathbin{\llcorner}B_{\mathbb{H}^{\mathfrak{m}}}(P,2)\bigr)+C(\mathfrak{m})$
$\leqslant e^{-\frac{1}{c_{0}\delta}}e^{2n}\cdot\bigl(x^{\mathfrak{m}}(P)\bigr)^{-n}M_{\mathbb{R}^{\mathfrak{m}}}(T)+C(\mathfrak{m})$
@@
where $C$ in (3.31) depends on $\mathfrak{m}$ and $\Gamma$. ∎
As the proof of Theorem 3.3 is essentially local, we in fact establish the following lemma.
@@
Let $\Gamma$ be a $C^{1}$ submanifold of the hyperplane at infinity with (1.8) holds, let $\rho_{\Gamma}$ be defined as in Section 2 and let $P\in G_{r}$ with $r<\tfrac{1}{2}\rho_{\Gamma}$. Assume that $T$ is an area-minimizing locally rectifiable $n$-current in $\mathbb{H}^{\mathfrak{m}}$ that is asymptotic to $\Gamma$, satisfying
$\partial T\mathbin{\llcorner}B_{\mathbb{H}^{\mathfrak{m}}}(P,2)=0,$ (3.32)
and (3.15) with $\delta=\bigl(e^{2}x^{\mathfrak{m}}(P)\bigr)^{1+\alpha}$. If in addition, for some small constant $c_{0}=c_{0}(\mathfrak{m},\Gamma)>0$, the bound
@@
By Theorem 3.3, for every $P\in G_{r}$ with $r<\tfrac{1}{2}\rho_{\Gamma}$, we have
$M_{\mathbb{H}^{\mathfrak{m}}}\bigl(T\mathbin{\llcorner}B_{\mathbb{H}^{\mathfrak{m}}}(P,1)\bigr)<C(\mathfrak{m},\Gamma)\bigl(M_{\mathbb{R}^{\mathfrak{m}}}(T)+1\bigr).$ (3.35)
@@
$\|\textbf{u}\|_{C^{1,1}(\bar{\Omega})}\leqslant\varepsilon\bigl(x^{\mathfrak{m}}(P),\mathfrak{m},\Gamma\bigr),$ (3.37)
where $\Omega$ is the Euclidean projection of $B_{\mathbb{H}^{\mathfrak{m}}}(\Phi(P),1)$ onto $H^{+}$, and
@@
Theorem 1.4 follows by the same argument, as it is essentially local.
@@ 4. System of Minimal Surface Equations
Setting $y^{1}=x^{1},\cdots,y^{n-1}=x^{n-1},y^{n}=x^{\mathfrak{m}}$ as in (1.12), we may rewrite this graph mapping as
@@
For convenience, we extend the coordinate $y$ on $H^{+}$ to a system of Euclidean coordinates in $\mathbb{R}^{\mathfrak{m}}$, where
$y^{n+1}=x^{n},\quad y^{n+2}=x^{n+1},\quad\cdots,\quad y^{\mathfrak{m}}=x^{\mathfrak{m}-1}.$ (4.7)
Then, for indices satisfying $1\leqslant i\leqslant n$ and $n+1\leqslant l\leqslant\mathfrak{m}$, we have
$\mathbf{F}_{i}=y^{i},\quad\mathbf{F}_{l}=\mathbf{u}_{l}.$ (4.8)
@@
If all mean curvatures $H_{s}$ vanish, then $\Delta_{g_{M}}\mathbf{F}=0$. To establish this, we start with the covariant derivative expansion of $\mathbf{F}_{,ij}$: for $1\leqslant i,j\leqslant n,$
@@
$\nabla_{\frac{\partial}{\partial y^{l}}}\frac{\partial}{\partial y^{k}}=-\frac{1}{y^{n}}\left(\delta_{ln}\frac{\partial}{\partial y^{k}}+\delta_{kn}\frac{\partial}{\partial y^{l}}-\delta_{lk}\frac{\partial}{\partial y^{n}}\right).$ (4.14)
@@
$g_{M}^{ij}\frac{\partial\mathbf{F}_{l}}{\partial y^{i}}\frac{\partial\mathbf{F}_{k}}{\partial y^{j}}\cdot\nabla_{\frac{\partial}{\partial y^{l}}}\frac{\partial}{\partial y^{k}}=\begin{cases}-\frac{1}{y^{n}}g_{M}^{lk}\left(\delta_{ln}\frac{\partial}{\partial y^{k}}+\delta_{kn}\frac{\partial}{\partial y^{l}}-\delta_{lk}\frac{\partial}{\partial y^{n}}\right),&1\leqslant l,k\leqslant n;\\ -\frac{1}{y^{n}}\frac{\partial u_{l-n}}{\partial y^{i}}\cdot\delta_{kn}g_{M}^{ik}\frac{\partial}{\partial y^{l}},&n+1\leqslant l\leqslant\mathfrak{m}\text{ and }1\leqslant k\leqslant n;\\ \frac{1}{y^{n}}\frac{\partial u_{l-n}}{\partial y^{i}}\frac{\partial u_{k-n}}{\partial y^{j}}\cdot\delta_{lk}g_{M}^{ij}\frac{\partial}{\partial y^{n}},&n+1\leqslant l,k\leqslant\mathfrak{m}.\end{cases}$ (4.15)
@@
If the graph of $\mathbf{u}\in C^{1}$ in the form given by (4.3), defined over a domain $\Omega\subseteq H^{+}$, is an $n$-dimensional minimal submanifold of $\mathbb{H}^{\mathfrak{m}}$ (i.e., all mean curvatures vanish), then the PDE system (4.17) holds.
We may also derive the system (4.17) via the method of variations. Consider an arbitrary smooth map defined on $G_{r}$, given in the form
@@ 5. Higher Order Regularity Theorems
This equation is uniformly elliptic after rescaling in a neighborhood of any point $P\in B^{+}_{r/2}$. Precisely, under the rescaled coordinate transformation